# Deep Learning LOS Line-of-Sight

### v3

@@
Adaptive 3D Geometry‐based Stochastic Channel Prediction for 3D DL Selection
@@ Abstract
This paper addresses the challenges of mobile user requirements in shadowing and multi-fading environments, focusing on the Downlink (DL) radio node selection based on Uplink (UL) channel estimation. One of the key issues tackled in this research is the prediction performance in scenarios where estimated channels are integrated. An adaptive deep learning approach is proposed to improve performance, offering a compelling alternative to traditional interpolation techniques for air-to-ground link selection on demand. Moreover, our study considers a 3D channel model, which provides a more realistic and accurate representation than 2D models, particularly in the context of 3D network node distributions. This consideration becomes crucial in addressing the complex multipath fading effects within geometric stochastic 3D 3GPP channel models in urban environments. Furthermore, our research emphasises the need for adaptive prediction mechanisms that carefully balance the trade-offs between DL link forecasted frequency response accuracy and the complexity requirements associated with estimation and prediction. This paper contributes to advancing 3D radio resource management by addressing these challenges, enabling more efficient and reliable communication for energy-constrained flying network nodes in dynamic environments.
@@
publicationid: pubid: Copyright  © 2023 Your Name
@@ I Introduction
Wireless communication networks are steadily gaining their popularity in all aspects of business and social life. Advances in wireless technologies, such as 5G and the upcoming 6G, promise to deliver faster data rates or reduced latency. Among others, these improvements will open up new use cases in the field of real-time applications, e. g. autonomous vehicles, wireless robotics control, and more. However, the critical challenge on the way to their realisation remains the reliability of the communication links. In outdoor scenarios, 2D terrestrial networks may be enhanced by the utilisation of the third coordinate by means of airborne wireless nodes.
3D radio resource nodes enable stable and highly reliable links to users, thanks to their high position flexibility within a 3D space. This capability addresses wireless signal propagation, blockage, reflection, shadowing, and the multipath effect. Compared to wireless links established through fixed terrestrial base stations, 3D radio resource nodes significantly improve wireless link quality.
@@
The flight node can serve as base stations or relays and has high flexibility to improve the connectivity link quality with optimal 3D trajectory planning or altitude change, unlike the fixed terrestrial radio access node with a low Line-of-Sight (LoS) probability [1]. In 3D networking, ongoing research explores the advantages of employing over-the-air network nodes to enhance Quality of Service (QoS) while minimising transmission time due to energy limitation.[2]. In the mobility control use case, through consecutive control input via a wireless channel, the radio access node selection can support a highly stable and reliable link considering the DL ground-to-ground link prediction with high accuracy and energy constraint of the flight nodes [3]. In  [4], the author optimised the wireless service to the ground user using optimal cell partitioning, considering the average hover times of the Unmanned Aerial Vehicles.
@@
Whereas utilising flight nodes is advantageous for maintaining a reliable communication link, the UAVs suffers from limited energy storage capability. To increase the energy efficiency of the flight nodes, we propose a 3D radio resource management system, which optimises the utilisation of power-demanding air-to-ground links based on QoS requirements. Thus, the ground-based users should preferably maintain the ground-to-ground link. In case of the ground link degradation due to i. e. fading, shadowing, or other effects, the air-to-ground link switch is performed preventing the violation of QoS requirements. Establishing an accurate 3D radio resource management system, relying on precise decision-making for DL adaptive transmission over-the-air links instead of ground-to-ground links is imperative [5]. So, the DL selection cannot be accurate when it is based on outdated channel information using uplink channel estimates due to the rapid variation in channel conditions, especially in dynamic environments such as urban areas with mobile User Equipment (UE).
Utilising Machine Learning (ML) methods for DL channel prediction enhances accuracy, mainly when the channel exhibits temporal correlation that influences the learning process of the predictor through interaction with channel state information. The channel predictor can capture the interaction between consecutive channel states in the time and frequency domain, thus benefiting from temporal channel correlation and enabling accurate channel prediction[6],[7]. Another important aspect is channel modelling, which incorporates 3D position instead of simplistic stochastic 2D channel modelling. A 3D channel model provides a more realistic perspective in the elevation angles (zenith) domain. This modelling approach included more parameter variables and their dependencies, which impact the channel variation [8]. [9] [10],[11].
To accomplish the task of optimal link selection, precise information on the condition of both ground-to-ground as well as air-to-ground links is required. In modern communication systems, the link condition is indicated by Mean Square Error (MSE). This measure is estimated based on the dedicated training signals sent in the UL channel. MSE is considered to be reciprocal, so the UL estimations is also utilised for DL communication. However, fast fading is known to destroy the reciprocity. Furthermore, especially real-time control applications may suffer the timeliness of the estimated MSE. The ability to predict DL channel behavior based on the UL estimation would significantly mitigate these effects.
This paper focuses on predicting Channel frequency responses (CFRs) in the context of 3D DL selection within an urban wireless signal propagation environment, utilising pilot-aided estimation methods and the 3D geometric stochastic 3GPP channel model considering the elevation domain. We analyse the prediction accuracy, considering the estimation error of the UL channel. Then, we introduce an innovative adaptive predictor structure for enhancing DL channel selection by leveraging upcoming channel information, thus ensuring mobile users’ highly reliable DL channel connection.
The paper is structured as follows. In the next section, the system model is presented. A detailed description of the simulation methodology follows in section III. Further, the results of our numerical analysis are discussed in detail in  IV. Finally, we summarise the main findings of this work in section V.
@@ II System model
In an Urban Microcell environment, mobile network status changes dynamically, necessitating DL selection for optimising 3D radio resources and ensuring radio link quality through a closed-loop continuous monitoring and optimisation process. Air-to-ground DL usage occurs on demand to conserve flight node radio resources, guided by decisions from the ground-to-ground DL selector. The proposed framework, illustrated in figure 1, deals with radio channel propagation problems considering the 3D channel modelling. The wireless channel between transmitter and mobile UEs in 3D dimension e.g. Automated Guided Vehicles or UAVs, exhibits both slow fading due to path loss and shadowing effects and fast fading owing to multipath effects. This consideration considers the 3D positions of both the receiver and transmitter, providing a highly accurate description of signal propagation in a 3D network.
@@ V Conclusion
We proceed to analyse the prediction gain between two prediction possibilities. The first option involves making predictions after each estimation in an open-loop fashion, while the second entails employing consecutive closed-loop predictions for improved accuracy. This choice represents a trade-off between estimation requirements and prediction accuracy. In case of estimation accuracy is high, consecutive closed-loop predictions can yield superior results. However, the closed-form prediction strategy may lose accuracy if the LS estimation accumulates estimation errors due to long-term dependencies.

### v5

@@
Adaptive Prediction Approach for 3D Geometry-based communication
@@ Abstract
This paper addresses the challenges of mobile user requirements in shadowing and multi-fading environments, focusing on the Downlink (DL) radio node selection based on Uplink (UL) channel estimation. One of the key issues tackled in this research is the prediction performance in scenarios where estimated channels are integrated. An adaptive deep learning approach is proposed to improve performance, offering a compelling alternative to traditional interpolation techniques for air-to-ground link selection on demand. Moreover, our study considers a 3D channel model, which provides a more realistic and accurate representation than 2D models, particularly in the context of 3D network node distributions. This consideration becomes crucial in addressing the complex multipath fading effects within geometric stochastic 3D 3GPP channel models in urban environments. Furthermore, our research emphasises the need for adaptive prediction mechanisms that carefully balance the trade-offs between DL forecasted frequency response accuracy and the complexity requirements associated with estimation and prediction. This paper contributes to advancing 3D radio resource management by addressing these challenges, enabling more efficient and reliable communication for energy-constrained flying network nodes in dynamic environments.
@@
IEEE© This work has been submitted to the IEEE conference for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Wireless communication networks are steadily gaining their popularity in all aspects of business and social life. Advances in wireless technologies, such as 5G and the upcoming 6G, promise to deliver faster data rates or reduced latency. Among others, these improvements will open up new use cases in the field of real-time applications, e. g: autonomous vehicles, wireless robotics control, and more. However, the critical challenge on the way to their realisation remains the reliability of the communication links. In outdoor scenarios, 2D terrestrial networks may be enhanced by the utilisation of the third coordinate using airborne wireless nodes. 3D radio resource nodes enable stable and highly reliable links to users, thanks to their high position flexibility within a 3D space. This capability addresses wireless signal propagation, blockage, reflection, shadowing, and the multipath effect. Compared to wireless links established through fixed terrestrial base stations, 3D radio resource nodes significantly improve wireless link quality. The flight nodes can serve as base stations or relays and has high flexibility to improve the connectivity link quality with optimal 3D trajectory planning or altitude change, unlike the fixed terrestrial radio access node with a low Line-of-Sight (LoS) probability [1]. In 3D networking, ongoing research explores the advantages of employing over-the-air network nodes to enhance Quality of Service (QoS) while minimising transmission time due to energy limitation [2]. In the mobility control use case, through consecutive control input via a wireless channel, the radio access node selection can support a highly stable and reliable link considering the ground-to-ground link prediction with high accuracy and energy constraint of the flight nodes [3]. The author optimised the wireless service to the ground user using optimal cell partitioning, considering the average hover times of the Unmanned Aerial Vehicles in [4].
@@
Whereas utilising flight nodes is advantageous for maintaining a reliable communication link, the UAVs suffers from limited energy storage capability. To increase the energy efficiency of the flight nodes, we propose a 3D radio resource management system, which optimises the utilisation of power-demanding air-to-ground links based on QoS requirements. Thus, the ground-based users should preferably maintain the ground-to-ground link. In case of the ground link degradation due to i. e. fading, shadowing, or other effects, the air-to-ground link selection is performed preventing the violation of QoS requirements. Establishing an accurate 3D radio resource management system, relying on precise decision-making for DL adaptive transmission over-the-air links instead of ground-to-ground links is imperative [5]. So, the DL selection cannot be accurate when it is based on outdated channel information using uplink channel estimates due to the rapid variation in channel conditions, especially in dynamic environments such as urban areas with mobile User Equipment (UE).
@@ II System model
In an Urban environment, channel state information (CSI) changes dynamically, necessitating DL selection for optimising 3D radio resources and ensuring radio link quality through a closed-loop continuous monitoring and optimisation process. Air-to-ground DL usage occurs on demand to save flight node radio resources and guarantee the provided user link reliability, guided by decisions from the ground-to-ground DL selector. The proposed framework, illustrated in figure 1, deals with radio channel propagation problems considering the 3D channel modelling. The wireless channel between transmitter and mobile UEs in 3D dimension, exhibits both slow fading and fast fading due to shadowing, multipath and doppler shift effects. This considers the 3D positions of both the receiver and transmitter and signal propagation in azimuth and elevation domain, providing a highly accurate description of signal propagation in a 3D network.
@@ V Conclusion
We proceed to analyse the prediction gain between two prediction possibilities. The first option involves making predictions after each estimation in an open-loop fashion, while the second entails employing consecutive closed-loop predictions for improved accuracy. This choice represents a trade-off between estimation requirements and prediction accuracy. In case estimation accuracy is high, consecutive closed-loop predictions can yield superior results. However, the closed-form prediction strategy may lose accuracy if the LS estimation accumulates estimation errors due to long-term dependencies.